Alpha Ltd is expecting annual earnings before interest and tax of ₹ 1.5 Lakhs. The company has 10% debentures of ₹ 4 lakhs and cost of Equity capital is 12%. Calculate the total value of the firm and the overall cost of capital of the firm according to Net Income Approach. Also comment what will happen to the value of the firm and the overall cost of capital if debt is increased in the capital structure.
1
Expert's answer
2020-06-11T10:56:24-0400

cost of capital:

EBIT=1.5

Less: Interest cost (10 %of 4)=0.4

Earnings (since tax is assumed to be absent)=1.1(1.5-0.4)

Shareholders’ Earnings=1.1Market value of Equity (1.1/12%)=9.17

Market value of Debt=4

Total Market value=13.17(4+9.17)

"cost of capital=\frac{EBIT}{Total value of firm}=1.5/13.17=0.1139 or 11.39%" %


total value of the firm=13.17(4+9.17)